OpenStudy (anonymous):

Jade made 12 3/4 cups of snack mix for his party. His guests ate 2/3 of the mix. How much snack mix did his guests eat. I need help ASAP please!!!

OpenStudy (anonymous):

12+(3/4) = (51/4) total snack mix (51/4)*(2/3) = (17/2) amount ate (51/4) - (17/2) = (17/4) = 4 and 1/4th left

OpenStudy (anonymous):

can you work this out slowly so i can follow you and understand please?

OpenStudy (anonymous):

You don't need the last part, since it isn't asking you how much is left

OpenStudy (anonymous):

12+(3/4) = (3/4) + (48/4) = (51/4)

OpenStudy (anonymous):

Does that make sense so far

OpenStudy (anonymous):

I understand the part where we multiply 12*4 and then add 3 we get 51/4 then from there i am lost

OpenStudy (anonymous):

Ok I am sorry I am still not understanding.

OpenStudy (anonymous):

Not good at math I think it would work if you went exactly step by step

OpenStudy (anonymous):

If you convert (51/4) into a decimal, you get: (2/3)rds of 12.75 = 8.5 8.5 = 17/2

OpenStudy (anonymous):

can you please go step by step telling me what to do first?

OpenStudy (anonymous):

http://www.mathsisfun.com/fractions_multiplication.html this should help you

OpenStudy (anonymous):

If you need to know how much of it they ate just divide the amount of mix into thirds and then double a third. since there are 3/4ths or literally 3 specific forths left you dont have to worry about that. 12 / 3 = what?

OpenStudy (anonymous):

Hi, can you tell me how to do this step by step from the beginning please larryboxplenty did not help very much lol!

OpenStudy (anonymous):

He helped quite alot you seem to be fishing for an exact answer and the steps to show it rather than help to find it

OpenStudy (anonymous):

I got the answer thank you, I would like an explanation. I am not fishing for answers that is why I am asking for steps

OpenStudy (anonymous):

To get the answer you basically need to figure out how much of that is 1/3. 3/4s divided by 3 is just a 4th so what ever you get just add a 1/4th on the end of it. 12 divided by 3 is what?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

this still doesn't help

OpenStudy (jdoe0001):

think of it this way if Jade had 4cups and her guests ate 1/2 of that, how much did they eat?

OpenStudy (anonymous):

3 1/2

OpenStudy (jdoe0001):

1/2 of 4 is 3 1/2?

OpenStudy (anonymous):

yeah I know that's what I said so then what?

OpenStudy (anonymous):

what is 4 divided by 2 ...

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Okay he asked you if they ate half of the 4..not a literal .5 of it

OpenStudy (anonymous):

I don't understand why mixed fractions are used since they are done away with in middle/high school, but meh :P Sorry I couldn't help, I thought you knew how to multiply fractions

OpenStudy (jdoe0001):

hmmm I see the confusion ...

OpenStudy (anonymous):

We know how to divide ok. I just need help to understand how to work out this problem.

OpenStudy (jdoe0001):

right, I meant 1/4 of 4, not 1/2 a cup of 4 cups

OpenStudy (jdoe0001):

well you see, the exercise is asking the same Jade started out with 12 3/4 "cups" and her guests ate 1/2 "amount" not 1/2 "cups"

OpenStudy (jdoe0001):

woops ahemm 2/3 rather well you see, the exercise is asking the same Jade started out with 12 3/4 "cups" and her guests ate 2/3 "amount" not 2/3 "cups"

OpenStudy (anonymous):

It wasn't 1/2 it was 2/3

OpenStudy (anonymous):

Ok so I have gotten as far as turning 12 3/4 into an improper fraction but do not know where to go from there.

OpenStudy (anonymous):

 This is 12 and (3/4). Not out of these, take away (2/3)rds as I will do in the next drawing

OpenStudy (anonymous):

 We have 3 groups, each with 17 parts